Suppose g(t)=t^2-4t.

a. Evaluate g(-3).

b. Solve g(t)=-3
Answer #1

g(t) = t2 - 4t

a) g(-3) = (-3)2 - 4(-3) = 9 + 12 = 21

b) g(t) = -3

=> t2 - 4t = -3

=> t2 - 4t + 3 = 0

=> t2 - 3t - t + 3 = 0

=> t(t - 3) - 1(t - 3) = 0

=> (t - 3)(t - 1) = 0

=> t = 3, t = 1

Therefore, the solutions are t = 3, t = 1